# An Agentic Framework with LLMs for Solving Complex Vehicle Routing Problems

### v1

@@ Appendix B Problem Statement
VRP Variant
Capacity
Open Route
Backhaul
Mixed Backhaul
Duration Limit
Time Windows
Multi depot
Electric Vehicle
@@
OVRP $\checkmark$ $\checkmark$
VRPB $\checkmark$ $\checkmark$
VRPL $\checkmark$ $\checkmark$
VRPTW $\checkmark$ $\checkmark$
OVRPTW $\checkmark$ $\checkmark$ $\checkmark$
OVRPB $\checkmark$ $\checkmark$ $\checkmark$
OVRPL $\checkmark$ $\checkmark$ $\checkmark$
VRPBL $\checkmark$ $\checkmark$ $\checkmark$
VRPBTW $\checkmark$ $\checkmark$ $\checkmark$
VRPLTW $\checkmark$ $\checkmark$ $\checkmark$
OVRPBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OVRPBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OVRPLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
VRPBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OVRPBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
VRPMB $\checkmark$ $\checkmark$
OVRPMB $\checkmark$ $\checkmark$ $\checkmark$
VRPMBL $\checkmark$ $\checkmark$ $\checkmark$
VRPMBTW $\checkmark$ $\checkmark$ $\checkmark$
OVRPMBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
VRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
MDOVRP $\checkmark$ $\checkmark$ $\checkmark$
MDVRPB $\checkmark$ $\checkmark$ $\checkmark$
MDVRPL $\checkmark$ $\checkmark$ $\checkmark$
MDVRPTW $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDVRPMB $\checkmark$ $\checkmark$ $\checkmark$
MDVRPMBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPMB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPMBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ C.3 Simulated Annealing Criterion
  • Note: Obj. denotes the objective value, and Gap(%) represents the relative gap with respect to HGS-PyVRP. Lower values indicate better performance. Bold numbers denote cases where the gap from SOTA is within 3%, which is considered acceptable for a fully automated and self-contained framework.
C.4 Comparison on 48 Standard Benchmark
The results of the additional 32 standard VRP variants are presented in Table 10, derived from 1,000 instances with 100 nodes each. AFL achieves results within 3% of the SOTA baseline HGS-PyVRP in most problem settings.
C.5 Details of Comparison of Code Reliability and Solution Feasibility
@@
C.6 Comparison on ATSP Benckmark
We evaluate our model on the ATSP benchmark (Johnson & McGeoch, 1997), which contains 18 instances with optimal solutions ranging from 17 to 443 nodes, to assess its applicability. The Asymmetric Traveling Salesman Problem (ATSP) is a variant of the classical TSP in which the distance from customer $i$ to customer $j$ may differ from the distance from customer $j$ to customer $i$, making the problem more challenging and representative of real-world routing scenarios such as one-way street networks or asymmetric transportation costs. The results, presented in Table 12, show that our model achieves consistently favorable performance on this task.
@@ C.6 Comparison on ATSP Benckmark
C.7 Comparison on ACVRP Benckmark

### v2

@@ Appendix B Problem Statement
VRP Variant Capacity (C) Open Route (O) Backhaul (B) Mixed Backhaul (MB) Duration Limit (L) Time Windows (TW) Multi depot (MD) Electric Vehicle (E)
@@
OCVRP $\checkmark$ $\checkmark$
CVRPB $\checkmark$ $\checkmark$
CVRPL $\checkmark$ $\checkmark$
CVRPTW $\checkmark$ $\checkmark$
OCVRPTW $\checkmark$ $\checkmark$ $\checkmark$
OCVRPB $\checkmark$ $\checkmark$ $\checkmark$
OCVRPL $\checkmark$ $\checkmark$ $\checkmark$
CVRPBL $\checkmark$ $\checkmark$ $\checkmark$
CVRPBTW $\checkmark$ $\checkmark$ $\checkmark$
CVRPLTW $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPMB $\checkmark$ $\checkmark$
OCVRPMB $\checkmark$ $\checkmark$ $\checkmark$
CVRPMBL $\checkmark$ $\checkmark$ $\checkmark$
CVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$
OCVRPMBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
MDOCVRP $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPB $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPL $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPTW $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPMB $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPMBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPMB $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPMBL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPMBTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPMBLTW $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ C.3 Simulated Annealing Criterion
Table 9: Gap comparison across different prompting strategies.
CVRP VRPL VRPTW OVRP VRPLTW OVRPL OVRPTW OVRPLTW
50 100 50 100 50 100 50 100 50 100 50 100 50 100 50 100
Self-Verification 7.61% 9.47% 6.80% 8.64% 5.71% 9.25% 3.54% 5.96%
COT 8.39% 12.04% 7.37% 10.51% 56.84% 93.83% 62.06% 105.86%
C.4 Code Generation Time Analysis
Table 10: Runtime Analysis Across Testing Phase and Solver Generation Phase.
Phase Component CVRP CVRPL CVRPTW OCVRP CVRPLTW OCVRPL OCVRPTW OCVRPLTW
Testing Problem Description 0.10 0.12 0.11 0.10 0.15 0.17 0.15 0.20
Solution Derivation 2.10 6.93 9.31 2.20 8.90 5.65 11.15 17.27
Solver Generation Problem Description 0.10 0.13 0.12 0.10 0.16 0.16 0.15 0.20
Code Generation 12.28 12.88 18.38 23.94 27.43 28.20 24.85 30.50
Solution Derivation 0.003 1.05 0.009 0.005 1.10 1.21 0.010 1.42
Figure 4: Comparison of solution quality across (a) different numbers of improvement iterations $T$ and (b) different testing time, evaluated over multiple VRP variants.
At first, we provide the runtime breakdown during testing phrase in Table 10. These results, as noted in Lines 350–351, measure the combined runtime for Problem Description and Solution Derivation over 1,000 test instances using the already generated codes.
A detailed runtime breakdown of solver generation phase for the three subtasks: Problem Description, Code Generation, and Solution Derivation, across eight representative VRP variants, are shown in Table 10. These results reflect the time required to generate executable solver code and obtain a feasible solution for a single instance, where the generated code can be reused in the same VRP variant. As shown in Table 10, the code can be generated within approximately 30 minutes, which is notably shorter than the time typically required for human expert design, highlighting the model’s strong efficiency in producing high-quality, executable, and feasible solver code within a short time.
In the Figure. 4, we explore the trade-off between solution quality and different number of solution refinement iterations.
@@ C.4 Code Generation Time Analysis
  • Note: Obj. denotes the objective value, and Gap(%) represents the relative gap w.r.t. HGS-PyVRP. Lower values indicate better performance. Bold numbers denote cases where the gap from SOTA is within 3%, which is considered acceptable for a fully automated and self-contained framework.
C.5 Comparison on 48 Standard Benchmark
The results of the all 48 standard VRP variants are presented in Table 11, derived from 1,000 instances with 100 nodes each. AFL achieves results within 3% of the SOTA baseline HGS-PyVRP in most problem settings.
C.6 Details of Comparison of Code Reliability and Solution Feasibility
@@
C.7 Comparison on ATSP Benckmark
We evaluate our model on the ATSP benchmark (Johnson and McGeoch, 1997), which contains 18 instances with optimal solutions ranging from 17 to 443 nodes, to assess its applicability. The Asymmetric Traveling Salesman Problem (ATSP) is a variant of the classical TSP in which the distance from customer $i$ to customer $j$ may differ from the distance from customer $j$ to customer $i$, making the problem more challenging and representative of real-world routing scenarios such as one-way street networks or asymmetric transportation costs. The results, presented in Table 13, show that our model achieves consistently favorable performance on this task.
@@ C.7 Comparison on ATSP Benckmark
C.8 Comparison on ACVRP Benckmark
@@ C.9 Comparison on SOP Benckmark
Table 16: Comparison results on different LLM.
n=50 n=100 n=50 n=100
HGS-PyVRP CVRP 10.37 10.40 15.62 20.80 CVRPL 10.59 10.40 15.77 20.80
GPT 4o 10.64 2.60 2.21 16.05 2.75 4.96 11.01 3.97 13.06 16.64 5.52 34.25
Claude 4 10.54 1.65 2.60 15.90 1.79 5.12 10.91 3.02 6.85 16.44 4.25 25.79
GPT 4.1 10.59 2.12 2.10 15.99 2.38 4.38 11.18 5.57 6.93 16.84 6.79 25.48
HGS-PyVRP CVRPTW 16.03 10.40 25.42 20.80 OCVRP 6.51 10.40 9.73 20.80
GPT 4o 16.45 2.62 12.62 26.88 5.74 25.44 6.87 5.53 3.07 10.36 6.47 7.38
Claude 4 16.28 1.56 8.50 26.19 3.03 33.78 6.60 1.38 2.95 9.94 2.26 7.22
GPT 4.1 16.19 0.99 9.31 25.79 1.46 38.45 6.64 2.00 2.20 9.99 2.67 5.52
HGS-PyVRP CVRPLTW 16.36 10.40 25.76 20.80 OCVRPL 6.51 10.40 9.72 20.80
GPT 4o 16.79 2.63 8.54 26.81 4.08 35.48 6.81 4.61 7.13 10.33 6.28 13.42
Claude 4 16.55 1.16 7.32 26.50 2.87 30.94 6.59 1.23 7.43 9.91 1.95 16.59
GPT 4.1 16.61 1.53 8.90 26.56 3.11 35.33 6.64 1.99 5.65 9.99 2.78 16.27
HGS-PyVRP OCVRPTW 10.51 10.40 16.93 20.80 OCVRPLTW 10.51 10.40 16.93 20.80
GPT 4o 10.91 3.81 11.89 17.68 4.43 27.39 10.88 3.52 11.29 17.66 4.31 58.89
Claude 4 10.63 1.14 14.89 17.24 1.83 35.95 10.64 1.24 12.23 17.26 1.95 62.24
GPT 4.1 10.55 0.38 11.15 17.04 0.65 49.05 10.58 0.67 17.27 17.19 1.54 70.31
C.10 AFL’s Sensitivity to Different LLMs
We tested AFL with two additional LLMs: Claude-Sonnet-4-20250514 and GPT-4o. The results are shown in Table 16. These experiments help us understand how different LLMs affect AFL’s performance and stability.
Solution Quality. From the results, we see that the strength of the LLM does influence solution quality. More powerful models generate more accurate and complete heuristics, which leads to better objective values. In our tests, the overall performance ranking is: Claude-Sonnet-4 >> GPT-4.1 >> GPT-4o. Although the final performance changes with different LLMs, AFL improves how well each model implements the required heuristics. As shown in Table 9, AFL performs better than both the standard prompt and alternative prompt strategies, showing that AFL can reliably boost a model’s ability to produce working heuristics.
Code Reliability. Most importantly, code reliability stays stable across all LLMs. No matter which model we use, AFL always generates executable code and feasible solutions. This shows that the multi-agent design keeps the whole pipeline stable even when the LLM is not very strong, proving that our framework is robust and widely applicable.
Table 17: Combined Results on JSON and CSV Format Inputs.
n=50 n=100 n=50 n=100
JSON CVRP 10.68 2.99 2.69 16.36 3.71 4.99 CVRPL 10.86 2.55 7.37 16.42 4.12 29.20
CSV 10.69 3.09 2.61 16.40 4.99 5.50 10.75 1.51 4.73 16.31 3.42 20.72
JSON CVRP TW 16.26 1.43 12.97 26.17 2.95 21.22 OCVRP 6.71 3.07 2.07 10.22 5.04 5.38
CSV 16.19 1.00 14.87 26.07 2.56 26.11 6.68 2.61 2.80 10.14 4.21 6.17
JSON CVRPL TW 16.63 1.65 8.02 26.52 2.95 31.64 OCVRP L 6.73 3.38 6.78 10.26 5.56 16.46
CSV 16.61 1.52 8.99 26.48 2.80 32.10 6.71 3.07 7.96 10.11 4.01 18.24
JSON OCVRP TW 10.58 0.67 14.73 17.19 1.54 55.47 OCVRP LTW 10.55 0.38 19.62 17.15 1.30 82.31
CSV 10.60 0.86 10.26 17.25 1.89 50.99 10.56 0.48 16.65 17.16 1.36 69.19
Table 18: Performance of AFL on CVRPLib-XXL
Gap(%) L1(3k) L2(4k) A1(6k) A2(7k) G1(10k) G2(11k) B1(15k) B2(16k)
POMO 75.30 78.16 112.27 159.22
LEHD 14.04 26.30 18.90 26.40 27.23 38.45 35.94 40.76
Our 8.12 13.34 8.60 13.62 8.02 14.85 7.01 16.56
C.11 Experiment on Different Input Format
As VRPLIB is the standard format in the VRP domain, we further evaluated AFL on JSON and CSV inputs to assess its robustness to alternative data representations commonly used in industrial pipelines and benchmark integrations. As shown in Table 17, the model consistently maintains strong performance and solution feasibility across different input formats.
C.12 Experiment on Large CVRP Instances
AFL is capable of handling extremely large VRP instances. Our current approach generates a general-purpose heuristic from a small instance (e.g., a 50-node instance, which can be obtained by trimming the data of the target instance and serves as a minimal example containing all relevant information such as constraints, problem characteristics, and format specifications). The generated solver is then applied to large-scale cases. As shown in Table 18, AFL achieves competitive performance on CVRPLib-XXL (Gao et al., 2024) with more than 16,000 customers.
During code generation, we also enforce a practical constraint: during solution derivation, the algorithm must execute within 10 seconds for the given instance. If this limit is exceeded, the system triggers a timeout, then the Revision Agent (RA) and Error Analysis Agent (EAA) are trigered to refine the algorithm for improved efficiency. We will add clarifications of this mechanism in the revised manuscript. Looking ahead, we will further enhance efficiency by exploring multi-objective evolutionary strategies to evolve the generated algorithms, jointly optimizing both performance and computational efficiency.